3. Solve: 323+1 = 243 (5 marks)
Answer #1

24+1 3 243 327+) : 243 Sol 243 - 35 pl22+5) 39 fafta) - 2011), then f(x)=3(x] 22+1 = 5 22- 5-) 2x=4 nale a 2:2